If I have 35.0 L of gas at room temperature (22.5o
C) and I compress it so that the
final volume is 12.0 L, what is the final temperature of the gas if the pressure is
constant?
1
Expert's answer
2020-05-01T14:41:54-0400

p1V1/T1=p2V2/T2;

p-const, so

V1/T1=V2/T2;

T2=T1*V2/V1=22.5*12/35=7.7=295.5*12/35=-171.68 oC